OpenStudy (anonymous):

Simplify the expression: [(14 + 2) • 4 – 1] • 6

OpenStudy (tkhunny):

Have you considered applying your Order of Operations?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

[(14+2)]=16 16*4=64 64+1=65 65*6=390 answer=390

OpenStudy (tkhunny):

...except that 64-1 = 63, not 65.

OpenStudy (anonymous):

i got it already, thanks anyways :) @tkhunny